OpenStudy (anonymous):

Rewrite with only sin x and cos x. cos 2x - sin x

sam (.sam.):

\[\cos (2x) - \sin x\] Use \[\cos(2x)=1-2\sin^2(x)\]

OpenStudy (anonymous):

andthen what?

OpenStudy (anonymous):

these are my answer choices a) cos2x - sin2x - sin x b) cos2x - sin3x c) cos2x + sin2x + sin x d)cos2x - 3 sin x

OpenStudy (anonymous):

b